# Achieving computational gains with quantum error correction primitives: Generation of long-range entanglement enhanced by error detection

### v1

@@ Abstract
The resource overhead required to achieve net computational benefits from quantum error correction (QEC) limits its utility while current systems remain constrained in size, despite exceptional progress in experimental demonstrations. In this paper, we demonstrate that the strategic application of QEC primitives without logical encoding can yield significant advantages on superconducting processors—relative to any alternative error-reduction strategy—while only requiring modest overhead. We first present a novel protocol for implementing long-range CNOT gates that relies on a unitarily-prepared Greenberger-Horne-Zeilinger (GHZ) state as well as a unitary disentangling step; the protocol natively introduces an error-detection process using the disentangled qubits as flags. We demonstrate that it achieves state-of-the-art gate fidelities of over 85% across up to 40 lattice sites, significantly and consistently outperforming the best alternative measurement-based protocol without introducing any additional ancilla qubits. We then apply sparse stabilizer measurements to generate large GHZ states by bit-flip and amplitude-damping errors. Employing this technique in combination with deterministic error suppression, we generate a 75-qubit GHZ state exhibiting genuine multipartite entanglement, the largest reported to date. The generation requires no more than 9 ancilla qubits and the fraction of samples discarded due to errors grows no higher than 78%, far lower than previous discard fractions required for tests using comparable numbers of fully encoded qubits. This work in total represents compelling evidence that adopting QEC primitives on current-generation devices can deliver substantial net benefits.
@@ I Introduction
Progress in experimental quantum error correction (QEC) has been rapid Krinner et al. (2022); Zhao et al. (2022); Acharya et al. (2023); Gupta et al. (2024); Bluvstein et al. (2024); Acharya et al. (2024); Reichardt et al. (2024); Matsos et al. (2024); Lachance-Quirion et al. (2023), with recent demonstrations achieving performance beyond break even, where QEC provides a net improvement in error rates over bare hardware. However, using QEC to achieve fault-tolerant computation still requires significant overheads Babbush et al. (2021), with estimates of up to 1000 physical qubits per logical qubit required to achieve error rates compatible with large-scale algorithms Dutkiewicz et al. (2024); Litinski (2019). In response, so-called early-fault-tolerance frameworks have been proposed Dutkiewicz et al. (2024), where overheads are reduced by targeting noise-resilient algorithms and achievement of less stringent logical-error rates. Even with these compromises, experimental demonstrations remain a challenge, and full QEC implementations on current-era devices typically deliver modest benefits at large computational costs in reduced qubit availability. Similarly, demonstrations of post-selected error-detection on experiments using a few dozen encoded logical qubits can have exceedingly high discard rates above 99.9% Bluvstein et al. (2024), posing another significant form of overhead for users with constrained access to hardware (by e.g., cost or time).
@@
In the interim, the community has developed a range of error-reduction techniques that do not require full QEC and its associated overheads. This begins with hardware-level innovation where progress in gate-error rates has been substantial Carroll et al. (2022); Bal et al. (2024); Norcia et al. (2024); Foss-Feig et al. (2024); Sete et al. (2024). To address the inevitable residual impact of noise encountered in even the most advanced devices, several quantum error mitigation approaches have been proposed as a near-term strategy for addressing readout Nation et al. (2021); Mundada et al. (2023) and computational errors Kandala et al. (2019); Huggins et al. (2021); van den Berg et al. (2023); Kim et al. (2023); Liao et al. (2023). These use post-processing of noise-corrupted quantum-computer outputs to improve the estimation of expectation values; this ensemble-post-processing approach is unfortunately largely incompatible with the fundamental concept of runtime error identification and correction underpinning QEC. In parallel, there has been rapid development of QEC-compatible routines based on gate and circuit-level error suppression Mundada et al. (2023); Coote et al. (2024); Carvalho et al. (2021); Tripathi et al. (2022); Sundaresan et al. (2020); Nation and Treinish (2023); Jones and Benjamin (2022); Hartnett et al. (2024); Pokharel et al. (2018); Seif et al. (2024). These approaches suppress errors in runtime and incur no overhead in circuit execution, enabling successful demonstrations of quantum algorithms at scales exceeding 100 qubits Sachdeva et al. (2024); Kim et al. (2023); Bluvstein et al. (2024); Liao et al. (2023). With current modestly sized devices, users generally achieve greater overall quantum computational capability by selecting these alternative strategies over QEC, in order to reduce error rates while preserving larger effective qubit counts. Nonetheless, these techniques cannot address all possible errors, with particular shortcomings in addressing (Markovian) incoherent computational errors such as pure dephasing or bit-flip errors.
@@
In this work, we demonstrate that net computational improvements can be achieved on current-generation quantum computers using low-overhead QEC primitives—the underlying protocols employed in QEC—without the need for full QEC encoding. We employ ancilla qubits and entangling parity checks (core elements of all stabilizer QEC codes Nielsen and Chuang (2011)) to perform error detection on unencoded computational qubits. We present two key demonstrations leveraging error-detection strategies to deliver enhanced overall capability. First, we improve the implementation of a long-range CNOT gate using a novel protocol based on unitary preparation of a Greenberger-Horne-Zeilinger (GHZ) state followed by a predominantly unitary disentangling step. This approach has the advantage that the final state of the disentangled qubits (the intermediate qubits that connect the distant control and target qubits themselves), reveals errors that have occurred during the gate. With this approach we perform a long-range CNOT achieving over 85% fidelity across up to 40 lattice sites, outperforming state-of-the-art alternatives on superconducting processors. Second, we generate large GHZ states using a protocol first presented by Mooney et al. Mooney et al. (2021), which allows for the integration of sparse error detection through ancillary stabilizer measurements. Using this resource-efficient routine (consuming no more than 9 flag qubits) in combination with a deterministic error-suppression pipeline Mundada et al. (2023); Coote et al. (2024); Hartnett et al. (2024), we achieve genuine multipartite entanglement for up to 75 qubits, verified in terms of multiple-quantum coherence (MQC) fidelity—a record in the published literature. In contrast to the work of Mooney et al., we see significant enhancements in the fidelity of the GHZ state through the addition of error detection with increasing numbers of flags. We also observe a comparatively low discard rate, where over $80\%$ of the shots are kept in the case of generating a 27-qubit GHZ state and over $21\%$ in the 75-qubit state. These results demonstrate that incorporating QEC primitives delivers a net improvement in the capability of a near-term quantum computer relative to the best alternative strategies, and (putting resource efficiency aside) achieves a much larger overall improvement than is currently achievable using full QEC.
@@ II.1 Novel circuit for long-range CNOTs
As a generic matter, long-range CNOTs can be implemented through a three-step framework, as depicted schematically in Fig. 1. First, one generates the shortest-path linear-chain GHZ state connecting the control and the target qubits. Second, one reduces the GHZ state to a Bell pair, with each member of the pair neighboring the control and the target qubits. In the third and final step, local operations and classical communication (LOCC) based on the Bell pair are used to implement the logical operation of the CNOT gate between the control and target qubits Eisert et al. (2000); Piveteau and Sutter (2024); Vazquez et al. (2024). By modifying the LOCC operations, this framework can be generalized to realize any controlled-unitary gate Eisert et al. (2000). The protocol described above is general; concrete implementations can vary significantly based on details of the GHZ-state-generation and Bell-pair-reduction steps.
Our new protocol follows this generalized three-step framework, and is depicted in Fig. 2(a) for the case of two qubits (Q0 and Q7) separated by six intermediate qubits (Q1–Q6). In step A, a linear-chain GHZ state is produced along the shortest path (Q1–Q6) between the target and control qubits, through repeated application of CNOT gates between adjacent qubits. The GHZ state is then reduced to a Bell pair on Q1 and Q6 through a combination of unitary operations and measurements in steps B and C. In particular, in step B, all intermediate GHZ data qubits except one are unitarily disentangled through the action of local CNOT gates, leaving behind a smaller GHZ state on Q1, Q3, and Q6. In step C, the last remaining GHZ data qubit Q3 is projectively disentangled through an $X$-basis measurement; its $X$-basis measurement outcome determines the parity of the remaining Bell pair on Q1 and Q6, hence the feedforward $Z$ gate. Other unitarily disentangled qubits (Q2, Q4, and Q5) can be optionally measured and used as flag qubits for error detection, which we explain in detail below. Lastly, step D proceeds as the LOCC step of the general framework discussed above to achieve the CNOT gate. In particular, we design both the generation and the reduction schemes (steps A to C) to rely primarily on unitary operations as opposed to projective measurements, better suiting the relative error budgets on current devices. We refer to our new protocol as the “unitary entangle-disentangle” protocol.
@@
Among the qubits in the GHZ state, the first and last are the neighbors of the control and target qubits, while the second lies at the halfway point of the teleportation path (rounded up or down by one lattice site if $n$ is even). This structure allows the disentangled qubits to serve as flag qubits; any deviation in readout on these $n-3$ qubits from the all-0 bit-string indicates the occurrence of an error. Bit-flip and amplitude damping errors that occur during the generation of the GHZ step or the reduction to Bell pair step may be captured, but due to the circuit structure, only $Z$-type errors that occur in the Hadamard gate in step A of the circuit in Fig. 2(a) are detectable.
Figure 3: Experimental results for long-range CNOT protocols. (a) The geometrical extent of the long-range CNOT, showing the control and target qubits (red) connected via a shortest path of intermediate qubits (green), which are prepared in a GHZ state during step A of the protocol. (b, c) Comparison of the average gate fidelity between the unitary entangle-disentangle (unitary ent.-disent.) protocol and the measurement-based protocol, as a function of teleportation distance (measured in qubits): (b) fidelity from shot-by-shot experiments and (c) fidelity after post-processing with additional readout error mitigation. Although both protocols feature conditional feedforward operations based on measurement outcomes, for simplicity in implementation, shots requiring feedforward were discarded, reducing the overall yield—though in all cases, a minimum remaining shot count of $1{,}000$ was ensured after overall post-selection. For the unitary entangle-disentangle protocol with error detection, erroneous shots due to the activation of flag qubit(s) were discarded. Data were collected consecutively on ibm_fez to ensure direct comparability without concern over hardware drift. Error bars represent ±$\pm 1$ standard deviation from $50$ bootstraps applied to a single experimental run.
@@ II.2 Experimental results
In Fig. 3(b) and (c) we present measurements of the effective gate fidelity as a function of system size. We determine the average gate fidelity using Monte Carlo process certification Flammia and Liu (2011); da Silva et al. (2011), following the procedure described in Ref. da Silva et al. (2011); Bäumer et al. (2024). The different panels present both a shot-by-shot mode, where each sample is considered in isolation, and a post-processing mode where multiple shots are aggregated statistically before the application of readout error mitigation as in Ref. Mundada et al. (2023), followed by post-selection. For each measurement approach, we separately present data with and without the use of error detection and compare against the state-of-the-art measurement-based protocol Bäumer et al. (2024). We do not experimentally implement the SWAP-based protocol due to its known inferior performance. All experiments incorporate error suppression in circuit execution to establish a high-fidelity baseline against which the addition of error detection can be evaluated in isolation  Mundada et al. (2023); Coote et al. (2024); Hartnett et al. (2024).
In both modes, the unitary entangle-disentangle protocol achieves consistently and significantly higher average gate fidelity than the measurement-based protocol for teleportation across up to at least $40$ qubits (Figs. 3(b) and (c)). We achieve $\gtrsim 90\%$ average gate fidelity up to 20Q in shot-by-shot experiments and up to 30Q when incorporating readout error mitigation in post-processing. Experiments with and without error detection both outperform the benchmark, demonstrating the relevance of the protocol design to performance. Taking advantage of the in-built error detection protocol reduces the effective gate infidelity by around $50\%$. The highest performance is achieved when also incorporating readout error mitigation in post-processing, consistent with the known error budgets for currently available devices, in which gate errors are lower than readout errors. For these experiments, the discarded fraction (fraction of experiments identified as failed in the error-detection routine) grows to approximately $80\%$ for the largest gate teleportation lengths in shot-by-shot mode: see App. B. There, we also provide an expanded comparison between the unitary entangle-disentangle protocol and the measurement-based protocol.
@@ III GHZ state preparation with error detection
GHZ states are maximally entangled multi-qubit states that carry significant roles in the broad field of quantum information. This includes use in entanglement-enhanced quantum sensing Degen et al. (2017), QEC, and multi-party quantum communications Hillery et al. (1999). The fidelity with which these states can be prepared in a quantum computer has also been proposed as a performance benchmark for quantum hardware Wei et al. (2020); Kam et al. (2024). These varied applications have led to substantial efforts attempting to generate large GHZ states on state-of-the-art quantum computers Mooney et al. (2021); Bäumer et al. (2024), involving both conventional unitary circuit and measurement-based preparation relying on mid-circuit measurement and feedforward (dynamic circuits) Andersen et al. (2020); Moses et al. (2023); Chen et al. (2023); Bäumer et al. (2024); Hashim et al. (2024).
In the computational basis, the $n$-qubit ($n>2$) GHZ state is represented
$|\text{GHZ}_{n}\rangle=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right).$ (2)
This makes clear the state’s extreme sensitivity to noise—a sensitivity which grows with the number of entangled qubits Monz et al. (2011). It is our objective to deploy QEC primitives in order to enhance the maximum circuit width capable of exhibiting genuine multipartite entanglement in the presence of realistic noise encountered in a quantum computer.
We adopt the strategy of unitary preparation and incorporation of sparse parity checks introduced by Mooney et al. Mooney et al. (2021), combining this technique with implementation of an error-suppression pipeline for arbitrary circuit execution Mundada et al. (2023). This general approach has the advantage of avoiding the penalties associated with readout errors in state preparation (see App. D for details on the inherent ineffectiveness of readout error mitigation for the measurement-based state-preparation strategy). As the implementation details of the error-suppression protocols have been published previously, we reserve discussion here to cover the relevant QEC primitives, and only include analysis of the impact of error suppression in our experimental measurements.
@@
In our implementation, a selected set of stabilizers of the form $Z_{i}Z_{j}$ is measured non-destructively by the use of a small number of additional ancilla qubits (flags). We classify the possible parity checks as type-$s$, where $s$ indicates the number of additional SWAP gates required for the error-detecting stabilizer measurement.

### v2

# Achieving computational gains with quantum error-correction primitives: Generation of long-range entanglement enhanced by error detection

@@ Abstract
The resource overhead required to achieve net computational benefits from quantum error correction (QEC) limits its utility while current systems remain constrained in size, despite exceptional progress in experimental demonstrations. In this paper, we demonstrate that the strategic application of QEC primitives without logical encoding can yield significant advantages on superconducting processors—relative to any alternative error-reduction strategy—while only requiring a modest overhead. We first present a novel protocol for implementing long-range CNOT gates that relies on a unitarily prepared Greenberger-Horne-Zeilinger (GHZ) state as well as a unitary disentangling step; the protocol natively introduces an error-detection process using the disentangled qubits as flags. We demonstrate that it achieves state-of-the-art gate fidelities of over 85% across up to 40 lattice sites, significantly and consistently outperforming the best alternative measurement-based protocol without introducing any additional ancilla qubits. We then apply sparse stabilizer measurements to generate large GHZ states by detecting bit-flip and amplitude-damping errors. Employing this technique in combination with deterministic error suppression, we generate a 75-qubit GHZ state exhibiting genuine multipartite entanglement, the largest reported to date. The generation requires no more than 9 ancilla qubits and the fraction of samples discarded due to errors grows no higher than 78%, far lower than previous discard fractions required for tests using comparable numbers of fully encoded qubits. This work in total represents compelling evidence that adopting QEC primitives on current-generation devices can deliver substantial net benefits.
@@ I Introduction
Progress in experimental quantum error correction (QEC) has been rapid Krinner et al. (2022); Zhao et al. (2022); Acharya et al. (2023); Gupta et al. (2024); Bluvstein et al. (2024); Acharya et al. (2025); Reichardt et al. (2024a); Matsos et al. (2024); Lachance-Quirion et al. (2023); Reichardt et al. (2024b), with recent demonstrations achieving performance beyond break even, where QEC provides a net improvement in error rates over bare hardware. However, using QEC to achieve fault-tolerant computation still requires significant overheads Babbush et al. (2021), with estimates of hundreds or thousands of physical qubits per logical qubit required to achieve error rates compatible with large-scale algorithms Litinski (2019). In response, so-called early-fault-tolerance frameworks have been proposed Dutkiewicz et al. (2024), where overheads are reduced by targeting noise-resilient algorithms and achievement of less stringent logical-error rates. Even with these compromises, experimental demonstrations remain a challenge, and full QEC implementations on current-era devices typically deliver modest benefits at high computational costs and a large overhead of physical qubits. Similarly, demonstrations of post-selected error detection on experiments using a few dozen encoded logical qubits can have exceedingly high discard rates above 99.9% Bluvstein et al. (2024), posing another significant form of overhead for users with constrained access to hardware (by e.g., cost or time).
@@
In the interim, the community has developed a range of error-reduction techniques that do not require full QEC and its associated overheads. This begins with hardware-level innovation where progress in qubit coherence and gate errors has been substantial Sundaresan et al. (2020); Carroll et al. (2022); Bal et al. (2024); Sete et al. (2024). To address the inevitable residual impact of noise encountered in even the most advanced devices, several quantum error mitigation approaches have been proposed as a near-term strategy for addressing readout Nation et al. (2021); Mundada et al. (2023) and computational errors Kandala et al. (2019); Huggins et al. (2021); van den Berg et al. (2023); Kim et al. (2023); Liao et al. (2024). These use post-processing of noise-corrupted quantum-computer outputs to improve the estimation of expectation values; this ensemble-post-processing approach is unfortunately largely incompatible with the fundamental concept of runtime error identification and correction underpinning QEC. In parallel, there has been rapid development of QEC-compatible routines based on gate and circuit-level error suppression Mundada et al. (2023); Coote et al. (2025); Seif et al. (2024); Tripathi et al. (2022); Pokharel et al. (2018); Carvalho et al. (2021); Nation and Treinish (2023); Hartnett et al. (2024). These approaches suppress errors at runtime and incur no overhead in circuit execution, enabling successful demonstrations of quantum algorithms at scales exceeding 100 qubits Sachdeva et al. (2024); Kim et al. (2023); Bluvstein et al. (2024); Liao et al. (2024). With current modestly sized devices, users generally achieve greater overall quantum computational capability by selecting these alternative strategies over QEC, in order to reduce error rates while preserving larger effective qubit counts. Nonetheless, these techniques cannot address all possible errors, with particular shortcomings in addressing (Markovian) incoherent computational errors such as pure dephasing or bit-flip errors.
@@
In this work, we demonstrate that net computational improvements can be achieved on current-generation quantum computers using low-overhead QEC primitives  [^1] By QEC primitives, we refer to core components or sub-routines employed in QEC, such as encoding, parity checks, non-destructive measurements via ancillas, and decoding.—the underlying protocols employed in QEC—on the physical level without the need for QEC encoding. We employ ancilla qubits and entangling parity checks (core elements of all stabilizer QEC codes Nielsen and Chuang (2011)) to perform error detection on unencoded computational qubits. We present two key demonstrations leveraging error-detection strategies to deliver enhanced overall capability. First, we improve the implementation of a long-range CNOT gate using a novel protocol based on unitary preparation of a Greenberger-Horne-Zeilinger (GHZ) state followed by a predominantly unitary disentangling step. This approach has the advantage that the final state of the disentangled qubits (the intermediate qubits that connect the distant control and target qubits themselves) reveals errors that have occurred during the gate. With this approach we perform a long-range CNOT achieving over 85% fidelity across up to 40 lattice sites, outperforming state-of-the-art alternatives on superconducting processors. Second, we generate large GHZ states using a protocol first presented by Mooney et al. Mooney et al. (2021), which allows for the integration of sparse error detection through ancillary stabilizer measurements. Using this resource-efficient routine (consuming no more than 9 flag qubits) in combination with a deterministic error-suppression pipeline Mundada et al. (2023); Coote et al. (2025), we achieve genuine multipartite entanglement for up to 75 qubits, verified in terms of multiple-quantum coherence (MQC) fidelity—a record in the published literature. In contrast to the work of Mooney et al., we see significant enhancements in the fidelity of the GHZ state through the addition of error detection with increasing numbers of flags. We also observe a comparatively low discard rate, where over $80\%$ of the shots are kept in the case of generating a 27-qubit GHZ state and over $21\%$ in the 75-qubit state. These results demonstrate that incorporating QEC primitives on the physical level can deliver a substantial net improvement in the capability of a near-term quantum computer relative to the best alternative strategies.
@@ II.1 Novel circuit for long-range CNOTs
As a generic matter, long-range CNOTs can be implemented through a three-step framework, as depicted schematically in Fig. 1. First, one generates the shortest-path linear-chain GHZ state connecting the control and the target qubits. In the computational basis, an $n$-qubit ($n>2$) GHZ state is represented
$|\psi_{\text{GHZ},n}\rangle=\frac{1}{\sqrt{2}}\left(\ket{0}^{\otimes n}+\ket{1}^{\otimes n}\right).$ (1)
Second, one reduces the GHZ state to a Bell pair, with each member of the pair neighboring the control and the target qubits. In the third and final step, local operations and classical communication (LOCC) based on the Bell pair are used to implement the logic of the CNOT gate between the control and target qubits Gottesman and Chuang (1999); Eisert et al. (2000); Piveteau and Sutter (2024); Vazquez et al. (2024). By modifying the LOCC operations, this framework can be generalized to realize any controlled-unitary gate Eisert et al. (2000). The protocol described above is general; concrete implementations can vary significantly based on details of the GHZ-state-generation and Bell-pair-reduction steps.
Our new protocol follows this generalized three-step framework, and is depicted in Fig. 2(a) for the case of two qubits (Q0 and Q7) separated by six intermediate qubits (Q1–Q6). In step A, a linear-chain GHZ state is produced along the shortest path (Q1–Q6) between the target and control qubits, through repeated application of CNOT gates between adjacent qubits. The GHZ state is then reduced to a Bell pair on Q1 and Q6 through a combination of unitary operations and measurements in steps B and C. In particular, in step B, all intermediate qubits (forming a GHZ state) except one are unitarily disentangled through the action of local CNOT gates, leaving behind a smaller GHZ state on Q1, Q3, and Q6. In step C, the last remaining GHZ data qubit Q3 is projectively disentangled through an $X$-basis measurement; its $X$-basis measurement outcome determines the parity of the remaining Bell pair on Q1 and Q6, hence the feedforward $Z$ gate. Other unitarily disentangled qubits (Q2, Q4, and Q5) can be optionally measured and used as flag qubits for error detection, which we explain in detail below. Lastly, step D proceeds as the LOCC step of the general framework discussed above to achieve the CNOT gate. In particular, we design both the generation and the reduction schemes (steps A to C) to rely primarily on unitary operations as opposed to projective measurements, better suiting the relative error budgets on current devices. We refer to our new protocol as the “unitary entangle-disentangle” protocol.
@@
Among the qubits in the GHZ state, the first and last are the neighbors of the control and target qubits, while the second lies at the halfway point of the teleportation path (rounded up or down by one lattice site if $n$ is even). This structure allows the disentangled qubits to serve as flag qubits; any deviation in readout on these $n-3$ qubits from the all-0 bit-string indicates the occurrence of an error. Bit-flip and amplitude damping errors that occur during the generation of the GHZ step or the reduction to Bell pair step may be captured, but due to the circuit structure, only $Z$-type errors that occur in the Hadamard gate in step A of the circuit in Fig. 2(a) are detectable. Detector theory McEwen et al. (2023) can also be applied here to identify the locations of the errors.
Figure 3: Experimental results for long-range CNOT protocols. (a) The geometrical extent of the long-range CNOT, showing the control and target qubits (red) connected via a shortest path of 25 intermediate qubits (green), which are prepared in a GHZ state during step A of the protocol. (b, c) Comparison of the average gate fidelity between the unitary entangle-disentangle (unitary ent.-disent.) protocol and the measurement-based protocol, as a function of teleportation distance (measured in qubits): (b) fidelity from shot-by-shot experiments and (c) fidelity after post-processing with additional readout error mitigation. Although both protocols feature conditional feedforward operations based on measurement outcomes, for simplicity in implementation, shots requiring feedforward were discarded, reducing the overall yield—though in all cases, a minimum remaining shot count of $1{,}000$ was ensured after overall post-selection. For the unitary entangle-disentangle protocol with error detection, erroneous shots due to the activation of flag qubit(s) were discarded. Data were collected consecutively on ibm_fez to ensure direct comparability without concern over hardware drift. Error bars represent ±$\pm 1$ standard deviation from $50$ bootstraps applied to a single experimental run.
@@ II.2 Experimental results
In Fig. 3(b) and (c) we present measurements of the effective gate fidelity as a function of system size. We determine the average gate fidelity using Monte Carlo process certification Flammia and Liu (2011); da Silva et al. (2011), following the procedure described in Ref. da Silva et al. (2011); Bäumer et al. (2024). The different panels present both a shot-by-shot mode, where each sample is considered in isolation, and a post-processing mode where multiple shots are aggregated statistically before the application of readout error mitigation as in Ref. Mundada et al. (2023) on all bits, followed by post-selection on the intermediate qubits. For each measurement approach, we separately present data with and without the use of error detection and compare against the state-of-the-art measurement-based protocol Bäumer et al. (2024). We do not experimentally implement the SWAP-based protocol due to its known inferior performance. All experiments incorporate error suppression in circuit execution Mundada et al. (2023); Coote et al. (2025) to establish a high-fidelity baseline against which the addition of error detection can be evaluated in isolation.
In both modes, the unitary entangle-disentangle protocol achieves consistently and significantly higher average gate fidelity than the measurement-based protocol for teleportation across up to at least $40$ qubits (Figs. 3(b) and (c)). We achieve $\gtrsim 90\%$ average gate fidelity up to 20Q in shot-by-shot experiments and up to 30Q when incorporating readout error mitigation in post-processing. Experiments with and without error detection both outperform the benchmark, demonstrating the relevance of the protocol design to performance. Taking advantage of the in-built error detection protocol reduces the effective gate infidelity by around $50\%$. The highest performance is achieved when also incorporating readout error mitigation in post-processing, consistent with the known error budgets for currently available devices, in which gate errors are lower than readout errors. It is important to emphasize the reasonably low runtime overhead due to the error detection of our protocol—for these experiments, the discarded fraction (fraction of experiments identified as failed in the error-detection routine) grows to at most approximately $80\%$ for the largest gate teleportation lengths in shot-by-shot mode: see App. B. There, we also provide an expanded comparison between the unitary entangle-disentangle protocol and the measurement-based protocol.
@@ III GHZ state preparation with error detection
GHZ states are maximally entangled multi-qubit states that carry significant roles in the broad field of quantum information. This includes use in entanglement-enhanced quantum sensing Degen et al. (2017), QEC, and multi-party quantum communications Hillery et al. (1999). The fidelity with which these states can be prepared in a quantum computer has also been proposed as a performance benchmark for quantum hardware Wei et al. (2020); Kam et al. (2024). These varied applications have led to substantial efforts attempting to generate large GHZ states on state-of-the-art quantum computers Mooney et al. (2021); Bäumer et al. (2024), involving both conventional unitary circuit and measurement-based preparation relying on mid-circuit measurement and feedforward (dynamic circuits) Andersen et al. (2020); Yang et al. (2022); Moses et al. (2023); Chen et al. (2025); Bäumer et al. (2024); Hashim et al. (2024).
The $n$-qubit GHZ state has an extreme sensitivity to noise—a sensitivity which grows with the number of entangled qubits Monz et al. (2011). It is our objective to deploy QEC primitives in order to enhance the maximum circuit width capable of exhibiting genuine multipartite entanglement in the presence of realistic noise encountered in a quantum computer.
We adopt the strategy of unitary preparation and incorporation of sparse parity checks introduced by Mooney et al. Mooney et al. (2021), combining this technique with implementation of an error-suppression pipeline for arbitrary circuit execution Mundada et al. (2023); Coote et al. (2025). This general approach has the advantage of avoiding the penalties associated with readout errors in state preparation (see App. D for details on the inherent ineffectiveness of readout error mitigation for the measurement-based state-preparation strategy). As the implementation details of the error-suppression protocols have been published previously, we reserve discussion here to cover the relevant QEC primitives, and only include analysis of the impact of error suppression in our experimental measurements.
@@
In our implementation, a selected set of stabilizers of the form $Z_{i}Z_{j}$ (each can be expressed as a product of some stabilizer generators defined in Eq. (3)) is measured non-destructively by the use of a small number of additional ancilla qubits (flags). We classify the possible parity checks as type-$s$, where $s$ indicates the number of additional SWAP gates required for the error-detecting stabilizer measurement.